OpenStudy (anonymous):

Find, by hand, three valid solutions to (d^2x/dt^2)^3 = tx(t),x(0)=x'(0)=0.

OpenStudy (anonymous):

Im not quite sure what to do with the first time, given that its a second order differential equation, but being to the 3rd power. I cant find examples like it. Should I be trying to sub in something for dx/dt?

OpenStudy (anonymous):

This is the eq. right?\[\left(\frac{d^2x}{dt^2}\right)^3=t~x(t)\]

OpenStudy (anonymous):

One way to do it would be with the Laplace transform. Have you covered that yet?

OpenStudy (anonymous):

Yes, sorry got busy with PuTTy.

OpenStudy (anonymous):

I haven't used Laplace with this style of problem though? Powers of derivatives that is

OpenStudy (anonymous):

Well all you need to know is how to find the transform of a derivative. This will give you a polynomial in terms of the transform of \(y\), which you can expand.

OpenStudy (anonymous):

L(f''(t)) = - f'(0) + sL(f'(t))

OpenStudy (anonymous):

Sorry, \(x\), not \(y\). If I remember it correctly, \[\mathcal{L}\{x''\}=s^2\mathcal{L}\{x\}-sx(0)-x'(0)=s^2\mathcal{L}\{x\}\] since \(x(0)=x'(0)=0\). \[\begin{align*} \left(s^2\mathcal{L}\{x\}\right)^3&=\mathcal{L}\{tx\}\\\\ s^6\left(\mathcal{L}\{x\}\right)^3&=-\mathcal{L}\{x\}\\\\ s^6\left(\mathcal{L}\{x\}\right)^3+\mathcal{L}\{x\}&=0\\\\ \mathcal{L}\{x\}\left(s^6\left(\mathcal{L}\{x\}\right)^2+1\right)&=0 \end{align*}\] This gives you two equations to solve.

OpenStudy (anonymous):

Ah yeah, I was looking at the wrong line. Sweet. Thanks for the help

OpenStudy (anonymous):

The first equation \(\mathcal{L}\{x\}=0\) gives the trivial solution. The second one might be a bit tricky to figure out, but at least it's quadratic.

OpenStudy (anonymous):

Yeah, thats manageable I believe

OpenStudy (anonymous):

It might also be worth noting that the Laplace transform of a product can be written as the convolution of their transforms: http://www.atp.ruhr-uni-bochum.de/rt1/syscontrol/node145.html#sec:AConvolutionInFrequencyDomain It's a bit above my knowledge of the integral transform, but it may be of interest to you.

OpenStudy (anonymous):

The other option would be to write guess solutions until you find something that works. You can eliminate a few functions, like the trigonometric. Something of the form \(e^{P(t)}\) might work, where \(P(t)\) is a polynomial with not-necessarily integral powers.

OpenStudy (anonymous):

It did tell me to try at^n

OpenStudy (anonymous):

It's nice when you get a hint. \[\begin{align*} x&=at^n\\ x'&=ant^{n-1}\\ x''&=an(n-1)t^{n-2}\\ (x'')^3&=a^3(n(n-1))^3t^{3n-6} \end{align*}\] \[\begin{align*} a^3(n(n-1))^3t^{3n-6}&=at^{n+1}\\ a^2(n(n-1))^3t^{2n-7}&=1\end{align*}\] The factor of \(t\) will have to disappear, and you can guarantee that by letting \(n=\dfrac{7}{2}\), since \(t^{2(7/2)-7}=t^0=1\) for \(t\not=0\) (I'm not yet sure if this will cause any problems with the solution...). You'll be able to find \(a\) pretty easily then.

OpenStudy (anonymous):

Sorry. I was gonna try to keep going myself, I just for the life of me couldnt think of how to start. Thanks again

OpenStudy (anonymous):

You're welcome